OpenStudy (anonymous):

Solve the equation 2 over 3 k + 6 = -12 for k. Please show your work

OpenStudy (anonymous):

\[\frac{2}{3k + 6} = -12 ?\]

OpenStudy (anonymous):

2 = -36k - 72 74 = -36k k = 74/-36

hero (hero):

moneybird, simplify it though

OpenStudy (anonymous):

37/-18

OpenStudy (anonymous):

2 over 3 k + 6 = -12 for k. (2/3)(k+6) = -12 multiply 2/3 and k+6 2k+12 over 3 = -12 to remove the denominator multiply both sides of the equation by 3 2k+12 = -36 to solve for k subtract 12 on both sides of the equation 2k+12-12 = -36-12 2k = - 48 divide both sides of the equation by 2 2k/2 = -48/2 k = -24